Agenda/Minutes on 2020/10/23
CRY+VIS Interface Meeting
Participants: Takahashi, Miyo, Ikeda, Hirata, Hagiwara, Aso, Matteo, Akutsu, Tomaru, Ushiba
- Schedule of safe mode switching for WAB: in 3 days
- Schedule of the wire disconnection: EY 11/3, IX and EX should be rearranged (Meeting on 11/27).
- Schedule of the FC painting.
IP monitoring improvement in vacuum pumping -> auto-tracking system: Test in the next venting on 11/4 and 11/20.
Participants: Takahashi, Miyo, Ikeda, Fabian, Hirata, Iguchi, Tanaka, Sugiyama, Terrence, Washimi, Aso
- We used Ikeda-san's and Miyo-kun's script for moving IPs with fishing rods during venting. It works very well and the interface is very good.
Set up the ribbon heaters in IX for the temperature control test. But the test conflicted with the vent of chamber. --> Move to EX and test them in the next week.
- Ikeda-san, Miyo-kun and Fabian monitored the IPs during venting.
- On Saturday Miyo-kun, Washimi-san and Fabian went to the tunnel and rerouted and reconnected cables in SR suspensions (IP and GAS LVDTs and FRs). We also reconnected the coil cables in BS OSEM satellite boxes for Guardian test.
- Hirata-san and I gave some feedback to Matteo about how we helped Hirose-san with the cleanliness of the mirrors before O3GK.
- In the context of Ikeda-san's work with stepper motor scripts, together with Miyo-kun I revisited the IP fishing rods transformation matrices.
- Before venting of the central room Fabian set the IPs and GAS filters in suitable positions (various klog reports). In the case of GAS filters it's useful to measure the keystone displacement due to bouyancy.
I checked SRM BF condition on Wednesday before the venting. See klog 15305. In previous analysis, it seemed the keystone hit a hard physical stop because it stopped moving when lowered. However, on Wednesday it didn't stop moving, but an oscillation at 40 mHz appeared. Currently I especulate it's a cable pulling.
Miyo-kun and I monitored the Type-B suspensions during venting. See klog 15307. I'll report the size of the bouyancy correction next week.
I wrote a To DO list for next week in Github: Type-B TODO list.
- Akutsu-san removed oplevs from SR suspensions as preparation to duct removal by Mirapro.
- As preparation for duct removal by Mirapro, we disconnected cables from PR2 small rack inside of the clean room and move it away from the chamber.
- (sato)No progress
- Design for the test board was finished. The board and parts will be ordered soon.
- The cost of two units for 3ch is about 600,000Yen. It is necessary to reduce the number of units in this FY.
- Chamber opening will be started from Oct. 26.
- Preparation work is on going.
- Monitoring for vent. Date: 19(EY), 21(IX, IY, EY), 22(IX, IY, EY, Center), 23(Center). Charge: Miyo, Ikeda, Fabian
- New guardian test on 19 in PR2 and BS.
- Procedure documents.
- Criteria in the acceptance test.
The week from 10/26
- Hirata: 10/26-30
- Takahashi: 10/26-28
The week from 11/2
- Takahashi: 11/1-6
- Sato: 11/1-6
- Hirata: 11/1-6
- On 2020/10/30(Fri)